OpenStudy (anonymous):

On Monday, David’s bank balance was $97. On Tuesday, his bank balance had changed by $56. Which amount could be the correct amount in David’s bank balance on Tuesday? A. $163 B. $56 C. $41 D. $9

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

It could've ADDED 56 to it or SUBTRACTED 56 to it. 97 - 56 = ? 97 + 56 = ?

OpenStudy (anonymous):

c?

OpenStudy (igreen):

Yep! 97 - 56 = 41 97 + 56 = 153 41 is the only one that's an option so that's the answer!